5/2 divided 5/7=
When I worked this problem I got 10/35 but I don't feel I divided correcty. help

Answer by checkley71(8403)

YOU ARE CORRECT --- YOU DIVIDED INCORRECTLY. YOU INVERTED THE NUMERATOR INSTEAD OF THE DENOMINATOR.
(5/2)/(5/7)= 5/2*7/5=35/10=7/2=3.5 ANSWER.
PROOF:
5/2=2.5
5/7=.7142857 NOW DIVIDE 2.5 BY .7142857
2.5/.7142857=3.5